# Distributed Safety Critical Control among Uncontrollable Agents using Reconstructed Control Barrier Functions

### 11 Mar 2026 v1

@@ Abstract
This paper investigates the distributed safety critical control for multi-agent systems (MASs) in the presence of uncontrollable agents with uncertain behaviors. To ensure system safety, the control barrier function (CBF) is employed in this paper. However, a key challenge is that the CBF constraints are coupled when MASs perform collaborative tasks, which depend on information from multiple agents and impede the design of a fully distributed safe control scheme. To overcome this, a novel reconstructed CBF approach is proposed. In this method, the coupled CBF is reconstructed by leveraging state estimates of other agents obtained from a distributed adaptive observer. Furthermore, a prescribed performance adaptive parameter is designed to modify this reconstruction, ensuring that satisfying the reconstructed CBF constraint is sufficient to meet the original coupled one. Based on the reconstructed CBF, we design a safety-critical quadratic programming (QP) controller and prove that the proposed distributed control scheme rigorously guarantees the safety of the MAS, even in the uncertain dynamic environments involving uncontrollable agents. The effectiveness of the proposed method is illustrated through simulations.
@@ I INTRODUCTION
In recent years, safety critical control has garnered significant attention for its applications in fields such as robotics [1], autonomous driving [2], and aerospace [3]. Beyond established methods such as artificial potential fields (APFs) [4], prescribed performance control (PPC) [5], and nonovershooting control [6], the use of control barrier functions (CBFs) to ensure system safety represents an emerging research direction. The standard CBF-based approach adopt a quadratic program (QP) as the controller to compute the control inputs that strictly satisfies the CBF-induced safety constraint, thereby guaranteeing system safety [7].
@@
Notation: Symbols $\mathbb{R}$ represents the real number set. $\mathbb{R}^{n}$ denotes the $n$-dimensional real vector space and $\mathbb{R}^{m\times n}$ denotes the $m\times n$-dimensional real matrix space. $I$ represents an identity matrix with an appropriate dimension. Inequalities $X\succ 0$ and $Y\succeq 0$ indicate that matrices $X$ and $Y$ are positive definite and positive semi-definite, respectively. $\mathbf{1}$ and $\mathbf{0}$ refer to vectors of proper dimensions with all entries to be 0 or $1$, respectively. The Kronecker product of matrices $M\in\mathbb{R}^{m\times n}$ and $N\in\mathbb{R}^{p\times q}$ is denoted by $M\otimes N\in\mathbb{R}^{mp\times nq}$.
@@ II-B System Dynamics and Control Barrier Functions
For a given ZCBF $h(x)$ and MAS (1), if the control input satisfy $u\in\mathcal{U}_{Z}$, where $\mathcal{U}_{Z}$ is the set of control input satisfying (4), the safe set $\mathcal{C}$ is forward invariant.
@@ II-C Problem Statement
Under Assumption 1, it is clearly that the non-zero diagonal matrix $\mathcal{B}_{j}\succeq 0$ for all $j\in\mathcal{N}^{+}$. Thus, it can be obtained that $\mathcal{H}_{j}\succ 0$, with the eigenvalues $\lambda_{j,l}$ of $\mathcal{H}_{j}$ satisfying $0<\lambda_{j,1}\leq\dots\leq\lambda_{j,N}$.
@@ III-A Design of Distributed Adaptive Observer
where $\tilde{f}_{i,l}:=f(\hat{x}_{i,l})-f(x_{i,l})$ and $d_{l}:=g({x}_{l})u_{l}$. Consequently, define $\tilde{x}_{l}:=\left[\tilde{x}_{1,l}^{T},\dots,\tilde{x}_{N,l}^{T}\right]^{T}$. Consider the following Lyapunov function
@@ III-B Reconstructed Control Barrier Functions
$\varepsilon_{i}=\frac{1}{2}\ln\left(\frac{e_{i}}{\rho_{i}-e_{i}}\right),\,\rho_{i}(t)=\left(\rho_{0}^{i}-\rho_{\infty}^{i}\right)\text{e}^{-\varrho t}+\rho_{\infty}^{i}.$
@@
1) Using proof by contradiction, we assume that $e_{i}$ first exceeds the performance bounds $(0,\rho_{i})$ at some finite time $\bar{t}$, which indicates that $\lim_{t\to\bar{t}}\varepsilon_{i}(t)\rightarrow\infty$. Hence, for $t\in[0,\bar{t})$, we have $e_{i}\in(0,\rho_{i})$. The following Lyapunov function is selected as
@@
$\left.\times\left(\dot{\vartheta}_{i}-\frac{dh(\hat{\upsilon}_{i})}{d\hat{\upsilon}_{i}}\dot{\hat{\upsilon}}_{i}+\frac{dh(x)}{dx}\dot{x}\right)\right]-\frac{1}{\gamma_{i}}\tilde{r}_{i}\dot{\hat{r}}$
@@
$+\frac{\varepsilon_{i}\rho_{i}}{2e_{i}\left(\rho_{i}-e_{i}\right)}\left[\frac{dh(x)}{dx}\left(\dot{x}_{i}-\dot{\hat{\upsilon}}_{i}\right)\right$.
$\left.+\left(\frac{dh(x)}{dx}-\frac{dh(\hat{\upsilon}_{i})}{d\hat{\upsilon}_{i}}\right)\dot{\hat{\upsilon}}_{i}\right]-\frac{1}{\gamma_{i}}\tilde{r}_{i}\dot{\hat{r}}$. (21)
@@
$+\frac{|\varepsilon_{i}|\rho_{i}}{2e_{i}\left(\rho_{i}-e_{i}\right)}\left[\left(\left\lVert\frac{dh(x)}{dx}\right\rVert+\left\lVert\dot{\hat{\upsilon}}_{i}\right\rVert\right)r_{i}\right]-\frac{1}{\gamma_{i}}\tilde{r}_{i}\dot{\hat{r}}$
@@
$\times\left[\left(\left\lVert\frac{dh(x)}{dx}\right\rVert+\left\lVert\dot{\hat{\upsilon}}_{i}\right\rVert-\left\lVert\frac{dh(\hat{\upsilon}_{i})}{d\hat{\upsilon}_{i}}\right\rVert-\left\lVert\dot{\hat{\bar{x}}}_{i}\right\rVert\right)r_{i}\right]$
@@
$+\frac{\chi_{i}^{2}\hat{r}_{i}^{2}}{\sqrt{\chi_{i}^{2}\hat{r}_{i}^{2}+\epsilon_{i}^{2}}}+\epsilon_{i}-\frac{1}{\gamma_{i}}\tilde{r}_{i}\dot{\hat{r}}$. (24)
@@
$\frac{|\varepsilon_{i}|\rho_{i}}{2e_{i}\left(\rho_{i}-e_{i}\right)}\left[\left(\left\lVert\frac{dh(x)}{dx}\right\rVert+\left\lVert\dot{\hat{\upsilon}}_{i}\right\rVert-\left\lVert\frac{dh(\hat{\upsilon}_{i})}{d\hat{\upsilon}_{i}}\right\rVert-\left\lVert\dot{\hat{\bar{x}}}_{i}\right\rVert\right)r_{i}\right]$
$\leq\frac{|\varepsilon_{i}|\rho_{i}}{2e_{i}\left(\rho_{i}-e_{i}\right)}\left[\left(\left\lVert\frac{dh(x)}{dx}-\frac{dh(\hat{\upsilon}_{i})}{d\hat{\upsilon}_{i}}\right\rVert+\left\lVert\dot{\hat{\upsilon}}_{i}-\dot{\hat{\bar{x}}}_{i}\right\rVert\right)r_{i}\right]$
@@
$\dot{V}_{i}\leq$ $-\frac{c_{i}}{2}\varepsilon_{i}^{2}+\varsigma_{i}\tilde{r}_{i}\hat{r}_{i}+\frac{1}{2}\Delta_{i}^{2}+\epsilon_{i}$
$\leq$ $-\frac{c_{i}}{2}\varepsilon_{i}^{2}-\frac{\varsigma_{i}}{2}\tilde{r}_{i}^{2}+\frac{\varsigma_{i}}{2}{r}_{i}^{2}+\frac{1}{2}\Delta_{i}^{2}+\epsilon_{i}$
@@
where $\zeta_{i}=\min\{c_{i},\varsigma_{i}\gamma_{i}\}$ and $\Omega_{i}:=\frac{\varsigma_{i}}{2}{r}_{i}^{2}+\frac{1}{2}\Delta_{i}^{2}+\epsilon_{i}$ is a positive constant. Integrating both sides of (III-B), it is obtained that
@@
From (27), it can be known that $V_{i}(\bar{t}^{-})$ is bounded, i.e., $\varepsilon_{i}(\bar{t}^{-})$ is bounded, which is in contradiction with the assumption $\lim_{t\to\bar{t}}\varepsilon_{i}(t)\rightarrow\infty$. Therefore, for $t\in[0,+\infty)$, $\varepsilon_{i}$ and $\tilde{r}_{i}$ remain bounded, i.e., reconstruction error $e_{i}$ satisfy the prescribed performance criteria.
@@ III-C Design of Distributed Safety Controller
Considering the MAS modeled by (1) under Assumption 1-5. With the proposed distributed control scheme (III-A), (III-B) and (28), the safe set $\mathcal{C}$ is forward invariant, i.e., the safety of all agents in $\mathcal{N}$ is guaranteed.
@@ IV SIMULATION
The robots’ working environment contains three circular obstacles $O_{i}$, $i=1,\dots,3$. The center position ${\boldsymbol{p}}_{O,i}$ and the radius $r_{O,i}$ of $O_{i}$ is set as: ${\boldsymbol{p}}_{O,1}^{T},r_{O,1}]=[0.8,2.5,0.5]$, ${\boldsymbol{p}}_{O,2}^{T},r_{O,2}]=[3,3.5,0.3]$, ${\boldsymbol{p}}_{O,3}^{T},r_{O,3}]=[3,1.5,0.5]$. Besides, the environment includes a goal point $G_{0}$ located at $\boldsymbol{p}_{G,0}=[0,4]^{T}$ and a circular target region $G_{1}$ centered at $\boldsymbol{p}_{G,1}=[1,0.5]^{T}$ with a radius $r_{G,1}=0.4$. Robot $4$ is required to navigate to the goal point $G_{0}$ and robots $1,2$ are tasked with navigating to the goal region $G_{1}$. In addition, robots $i$ for $i=1,2,3$ are subject to the following constraints:
@@
$h_{i}=-\frac{1}{20}\ln\left(\sum_{l=1}^{3}\text{e}^{-20b_{i,O_{l}}}+\sum_{j\neq i,j\in\mathcal{N}^{+}}\text{e}^{-20b_{i,j}}+\text{e}^{-20b_{i}}\right).$
@@
The initial states of the robots are set as follows: $\boldsymbol{p}_{1}(0)^{T},\theta_{1}(0)]=[4,4,0]$, $\boldsymbol{p}_{2}(0)^{T},\theta_{2}(0)]=[4,3,0]$, $\boldsymbol{p}_{3}(0)^{T},\theta_{3}(0)]=[4,0.5,-\pi]$, $\boldsymbol{p}_{4}(0)^{T},\theta_{4}(0)]=[3.5,0.6,-\pi]$. The distributed observers here are used solely for estimating $\boldsymbol{p}_{i}$ and the distributed observers parameters are set as: the initial estimate $\hat{x}_{i,j}$ is set to be identical to the actual initial state, $\hat{\delta}_{i,j}(0)=2$, $\sigma_{j}=0.01$, $P_{j}=\mathrm{diag}\{2,2\}$ for $i=1,2,3,j\in\mathcal{N}^{+}$. The parameters for the reconstructed CBFs are set as follows: $\theta_{i}(0)=0.1$, $\hat{r}_{i}(0)=0$, $\rho_{0}^{i}=1$, $\rho_{\infty}^{i}=0.15$, $\varrho=1$, $c_{i}=0.01$, $\varsigma_{i}=0.8$, $\gamma_{i}=0.01$, $\epsilon_{i}=0.01$ for $i=1,2,3$. For QP controller, we choose $\alpha_{1}(\hat{h}_{1})=\hat{h}_{1}$, $\alpha_{2}(\hat{h}_{2})=\hat{h}_{2}$, $\alpha_{3}(\hat{h}_{3})=\hat{h}_{3}^{5}/10$, and $W_{i}=g(\theta_{i})^{T}g(\theta_{i})$ for $i=1,2,3$. Additionally, the constraints on the robots’ maximum linear velocity and angular velocity are imposed, such that $|v_{i}|\leq v_{\max}$ and $|\omega_{i}|\leq\omega_{\max}$ for $i\in\mathcal{N}^{+}$, where $v_{\max},\omega_{\max}]=[0.22,2.84]$.

### 20 Mar 2026 v2

@@ Abstract
This paper investigates the distributed safety critical control for multi-agent systems (MASs) in the presence of uncontrollable agents with uncertain behaviors. To ensure system safety, the control barrier function (CBF) is employed in this paper. However, a key challenge is that the CBF constraints are coupled when MASs perform collaborative tasks, which depend on information from multiple agents and impede the design of a fully distributed safe control scheme. To overcome this, a novel reconstructed CBF approach is proposed. In this method, the coupled CBF is reconstructed by leveraging state estimates of other agents obtained from a distributed adaptive observer. Furthermore, a prescribed performance adaptive parameter is designed to modify this reconstruction, ensuring that satisfying the reconstructed CBF constraint is sufficient to meet the original coupled one. Based on the reconstructed CBF, we design a safety-critical quadratic programming (QP) controller and prove that the proposed distributed control scheme rigorously guarantees the safety of the MAS, even in the uncertain dynamic environments involving uncontrollable agents. The effectiveness of the proposed method is illustrated through a simulation.
@@ I INTRODUCTION
In recent years, safety critical control has garnered significant attention for its applications in fields such as robotics [1], autonomous driving [2], and aerospace [3]. Beyond established methods such as artificial potential fields (APFs) [4], prescribed performance control (PPC) [5], and nonovershooting control [6], the use of control barrier functions (CBFs) to ensure system safety represents an emerging research direction. The standard CBF-based approach adopts a quadratic program (QP) as the controller to compute the control inputs that strictly satisfies the CBF-induced safety constraint, thereby guaranteeing system safety [7].
@@
Notation: Symbol $\mathbb{R}$ represents the real number set. $\mathbb{R}^{n}$ denotes the $n$-dimensional real vector space and $\mathbb{R}^{m\times n}$ denotes the $m\times n$-dimensional real matrix space. $I$ represents an identity matrix with an appropriate dimension. Inequalities $X\succ 0$ and $Y\succeq 0$ indicate that matrices $X$ and $Y$ are positive definite and positive semi-definite, respectively. $\mathbf{1}$ and $\mathbf{0}$ refer to vectors of proper dimensions with all entries to be $1$ or 0, respectively. The Kronecker product of matrices $M\in\mathbb{R}^{m\times n}$ and $N\in\mathbb{R}^{p\times q}$ is denoted by $M\otimes N\in\mathbb{R}^{mp\times nq}$.
@@ II-B System Dynamics and Control Barrier Functions
For a given ZCBF $h(x)$ and MAS (1), if the control input satisfies $u\in\mathcal{U}_{Z}$, where $\mathcal{U}_{Z}$ is the set of control input satisfying (4), the safe set $\mathcal{C}$ is forward invariant.
@@ II-C Problem Statement
Under Assumption 1, it is clear that the non-zero diagonal matrix $\mathcal{B}_{j}\succeq 0$ for all $j\in\mathcal{N}^{+}$. Thus, it can be obtained that $\mathcal{H}_{j}\succ 0$, with the eigenvalues $\lambda_{j,l}$ of $\mathcal{H}_{j}$ satisfying $0<\lambda_{j,1}\leq\dots\leq\lambda_{j,N}$.
@@ III-A Design of Distributed Adaptive Observer
where $\tilde{f}_{i,l}:=f(\hat{x}_{i,l})-f(x_{l})$ and $d_{l}:=g({x}_{l})u_{l}$. Consequently, define $\tilde{x}_{l}:=\left[\tilde{x}_{1,l}^{T},\dots,\tilde{x}_{N,l}^{T}\right]^{T}$. Consider the following Lyapunov function
@@ III-B Reconstructed Control Barrier Functions
$\varepsilon_{i}=\frac{1}{2}\ln\left(\frac{e_{i}}{\rho_{i}-e_{i}}\right),\,\rho_{i}(t)=\left(\rho_{0}^{i}-\rho_{\infty}^{i}\right)\mathrm{e}^{-\varrho t}+\rho_{\infty}^{i}.$
@@
1) Using proof by contradiction, we assume that $e_{i}$ first exceeds the performance bounds $(0,\rho_{i})$ at some finite time $\bar{t}$, which indicates that $\lim_{t\to\bar{t}}\varepsilon_{i}(t)=\infty$. Hence, for $t\in[0,\bar{t})$, we have $e_{i}\in(0,\rho_{i})$. The following Lyapunov function is selected as
@@
$\left.\times\left(\dot{\vartheta}_{i}-\frac{dh(\hat{\upsilon}_{i})}{d\hat{\upsilon}_{i}}\dot{\hat{\upsilon}}_{i}+\frac{dh(x)}{dx}\dot{x}\right)\right]-\frac{1}{\gamma_{i}}\tilde{r}_{i}\dot{\hat{r}}_{i}$
@@
$+\frac{\varepsilon_{i}\rho_{i}}{2e_{i}\left(\rho_{i}-e_{i}\right)}\left[\frac{dh(x)}{dx}\left(\dot{x}-\dot{\hat{\upsilon}}_{i}\right)\right$.
$\left.+\left(\frac{dh(x)}{dx}-\frac{dh(\hat{\upsilon}_{i})}{d\hat{\upsilon}_{i}}\right)\dot{\hat{\upsilon}}_{i}\right]-\frac{1}{\gamma_{i}}\tilde{r}_{i}\dot{\hat{r}}_{i}.$ (21)
@@
$+\frac{|\varepsilon_{i}|\rho_{i}}{2e_{i}\left(\rho_{i}-e_{i}\right)}\left(\left\lVert\frac{dh(x)}{dx}\right\rVert+\left\lVert\dot{\hat{\upsilon}}_{i}\right\rVert\right)r_{i}-\frac{1}{\gamma_{i}}\tilde{r}_{i}\dot{\hat{r}}$
@@
$\times\left(\left\lVert\frac{dh(x)}{dx}\right\rVert+\left\lVert\dot{\hat{\upsilon}}_{i}\right\rVert-\left\lVert\frac{dh(\hat{\upsilon}_{i})}{d\hat{\upsilon}_{i}}\right\rVert-\left\lVert\dot{\hat{\bar{x}}}_{i}\right\rVert\right)r_{i}$
@@
$+\frac{\chi_{i}^{2}\hat{r}_{i}^{2}}{\sqrt{\chi_{i}^{2}\hat{r}_{i}^{2}+\epsilon_{i}^{2}}}+\epsilon_{i}-\frac{1}{\gamma_{i}}\tilde{r}_{i}\dot{\hat{r}}_{i}.$ (24)
@@
$\frac{|\varepsilon_{i}|\rho_{i}}{2e_{i}\left(\rho_{i}-e_{i}\right)}\left(\left\lVert\frac{dh(x)}{dx}\right\rVert+\left\lVert\dot{\hat{\upsilon}}_{i}\right\rVert-\left\lVert\frac{dh(\hat{\upsilon}_{i})}{d\hat{\upsilon}_{i}}\right\rVert-\left\lVert\dot{\hat{\bar{x}}}_{i}\right\rVert\right)r_{i}$
$\leq\frac{|\varepsilon_{i}|\rho_{i}}{2e_{i}\left(\rho_{i}-e_{i}\right)}\left(\left\lVert\frac{dh(x)}{dx}-\frac{dh(\hat{\upsilon}_{i})}{d\hat{\upsilon}_{i}}\right\rVert+\left\lVert\dot{\hat{\upsilon}}_{i}-\dot{\hat{\bar{x}}}_{i}\right\rVert\right)r_{i}$
@@
$\dot{V}_{i}\leq$ $-\frac{c_{i}}{2}\varepsilon_{i}^{2}+\frac{\varsigma_{i}}{\gamma_{i}}\tilde{r}_{i}\hat{r}_{i}+\frac{1}{2}\Delta_{i}^{2}+\epsilon_{i}$
$\leq$ $-\frac{c_{i}}{2}\varepsilon_{i}^{2}-\frac{\varsigma_{i}}{2\gamma_{i}}\tilde{r}_{i}^{2}+\frac{\varsigma_{i}}{2\gamma_{i}}{r}_{i}^{2}+\frac{1}{2}\Delta_{i}^{2}+\epsilon_{i}$
@@
where $\zeta_{i}=\min\{c_{i},\varsigma_{i}\}$ and $\Omega_{i}:=\frac{\varsigma_{i}}{2\gamma_{i}}{r}_{i}^{2}+\frac{1}{2}\Delta_{i}^{2}+\epsilon_{i}$ is a positive constant. Integrating both sides of (III-B), it is obtained that
@@
From (27), it can be known that $V_{i}(\bar{t}^{-})$ is bounded, i.e., $\varepsilon_{i}(\bar{t}^{-})$ is bounded, which is in contradiction with the assumption $\lim_{t\to\bar{t}}\varepsilon_{i}(t)=\infty$. Therefore, for $t\in[0,+\infty)$, $\varepsilon_{i}$ and $\tilde{r}_{i}$ remain bounded, i.e., reconstruction error $e_{i}$ satisfy the prescribed performance criteria.
@@ III-C Design of Distributed Safety Controller
Consider the MAS modeled by (1) under Assumption 1-5. With the proposed distributed control scheme (III-A), (III-B) and (28), the safe set $\mathcal{C}$ is forward invariant, i.e., the safety of all agents in $\mathcal{N}$ is guaranteed.
@@ IV SIMULATION
The robots’ working environment contains three circular obstacles $O_{i}$, $i=1,\dots,3$. The center position ${\boldsymbol{p}}_{O,i}$ and the radius $r_{O,i}$ of $O_{i}$ are set as: ${\boldsymbol{p}}_{O,1}^{T},r_{O,1}]=[0.8,2.5,0.5]$, ${\boldsymbol{p}}_{O,2}^{T},r_{O,2}]=[3,3.5,0.3]$, ${\boldsymbol{p}}_{O,3}^{T},r_{O,3}]=[3,1.5,0.5]$. Besides, the environment includes a goal point $G_{0}$ located at $\boldsymbol{p}_{G,0}=[0,4]^{T}$ and a circular target region $G_{1}$ centered at $\boldsymbol{p}_{G,1}=[1,0.5]^{T}$ with a radius $r_{G,1}=0.4$. Robot $4$ is required to navigate to the goal point $G_{0}$ and robots $1,2$ are tasked with navigating to the goal region $G_{1}$. In addition, robots $i$ for $i=1,2,3$ are subject to the following constraints:
@@
$h_{i}=-\frac{1}{20}\ln\left(\sum_{l=1}^{3}\mathrm{e}^{-20b_{i,O_{l}}}+\sum_{j\neq i,j\in\mathcal{N}^{+}}\mathrm{e}^{-20b_{i,j}}+\mathrm{e}^{-20b_{i}}\right).$
@@
The initial states of the robots are set as follows: $\boldsymbol{p}_{1}(0)^{T},\theta_{1}(0)]=[4,4,0]$, $\boldsymbol{p}_{2}(0)^{T},\theta_{2}(0)]=[4,3,0]$, $\boldsymbol{p}_{3}(0)^{T},\theta_{3}(0)]=[4,0.5,-\pi]$, $\boldsymbol{p}_{4}(0)^{T},\theta_{4}(0)]=[3.5,0.6,-\pi]$. The distributed observers here are used solely for estimating $\boldsymbol{p}_{i}$ and the distributed observers parameters are set as: the initial estimate $\hat{x}_{i,j}$ is set to be identical to the actual initial state, $\hat{\delta}_{i,j}(0)=2$, $\sigma_{j}=0.01$, $P_{j}=\mathrm{diag}\{2,2\}$ for $i=1,2,3,j\in\mathcal{N}^{+}$. The parameters for the reconstructed CBFs are set as follows: $\vartheta_{i}(0)=0.1$, $\hat{r}_{i}(0)=0$, $\rho_{0}^{i}=1$, $\rho_{\infty}^{i}=0.15$, $\varrho=1$, $c_{i}=0.01$, $\varsigma_{i}=0.8$, $\gamma_{i}=0.01$, $\epsilon_{i}=0.01$ for $i=1,2,3$. For QP controller, we choose $\alpha_{1}(\hat{h}_{1})=\hat{h}_{1}$, $\alpha_{2}(\hat{h}_{2})=\hat{h}_{2}$, $\alpha_{3}(\hat{h}_{3})=\hat{h}_{3}^{5}/10$, and $W_{i}=\mathrm{diag}\{1,l^{2}\}$ for $i=1,2,3$. Additionally, the constraints on the robots’ maximum linear velocity and angular velocity are imposed, such that $|v_{i}|\leq v_{\max}$ and $|\omega_{i}|\leq\omega_{\max}$ for $i\in\mathcal{N}^{+}$, where $v_{\max},\omega_{\max}]=[0.22,2.84]$.